# DAIEN-TTS: Disentangled Audio Infilling for Environment-Aware Text-to-Speech Synthesis

### v1

@@ Abstract
This paper presents DAIEN-TTS, a zero-shot text-to-speech (TTS) framework that enables ENvironment-aware synthesis through Disentangled Audio Infilling. By leveraging separate speaker and environment prompts, DAIEN-TTS allows independent control over the timbre and the background environment of the synthesized speech. Built upon F5-TTS, the proposed DAIEN-TTS first incorporates a pretrained speech-environment separation (SES) module to disentangle the environmental speech into mel-spectrograms of clean speech and environment audio. Two random span masks of varying lengths are then applied to both mel-spectrograms, which, together with the text embedding, serve as conditions for infilling the masked environmental mel-spectrogram, enabling the simultaneous continuation of personalized speech and time-varying environmental audio. To further enhance controllability during inference, we adopt dual class-free guidance (DCFG) for the speech and environment components and introduce a signal-to-noise ratio (SNR) adaptation strategy to align the synthesized speech with the environment prompt. Experimental results demonstrate that DAIEN-TTS generates environmental personalized speech with high naturalness, strong speaker similarity, and high environmental fidelity. Audio samples can be accessed at https://yxlu-0102.github.io/DAIEN-TTS.
@@ 2.3.2 SNR Adaption
Accordingly, we can derive that $Scale=\sqrt{\|\boldsymbol{Y}^{S}_{env}\|_{2}^{2}/\|\boldsymbol{Y}^{S}_{spk}\|_{2}^{2}}$. During inference, $\boldsymbol{Y}^{E}_{env}$ is scaled by the calculated factor and then transformed into a mel-spectrogram, which serves as the final environment condition for environmental speech generation.
@@ 3.1 Dataset and Experimental Setup
In the SES module, we set the number of Transformer layers $K$ to 8, with 16 attention heads, an embedding dimension of 1024, and a feed-forward network (FFN) dimension of 2048. In the TTS module, each DiT block is augmented with a cross-attention layer, also configured with 16 attention heads. During training, we adopted a dynamic mixing strategy to construct clean–environmental speech pairs. For pretraining the SES module, all speech data was mixed with environment audio to ensure robust separation learning. For TTS training, we applied a probabilistic augmentation strategy, where each speech sample was mixed with environmental audio with a probability of 50%, and with silent segments otherwise, in order to encourage better alignment learning between text and speech. Both the SES module and the TTS module were trained to 600k steps with a batch size of 102,800 audio frames on 24 NVIDIA V100 32G GPUs, following the same training strategy as F5-TTS. For inference, we set the DCFG strength $\alpha_{speech}$ and $\alpha_{env}$ to 2.0.
@@ 4.1 Evaluation with Silence Environment Prompts
The results of environment-aware TTS with silence environment prompts are shown in the upper part of Table 1. We first took the human recordings from the Seed-TTS test-en set as the ground truth (denoted as “Human”), and resynthesized them with the vocoder (denoted as “Vocoder”). Experimental results demonstrated that the vocoder resynthesized ground truth exhibited slight degradations in SIM-o and WER, which can be served as the upper bound for clean synthesized speech in terms of objective metrics. For the F5-TTS baseline trained on clean speech data, the synthesized speech exhibited high naturalness, intelligibility, and speaker similarity when provided clean speaker prompts. However, a noticeable drop in both objective and subjective metrics was observed when environmental speaker prompts were conditioned, indicating that F5-TTS was sensitive to background environment.

### v2

@@ Abstract
This paper presents DAIEN-TTS, a zero-shot text-to-speech (TTS) framework that enables ENvironment-aware synthesis through Disentangled Audio Infilling. By leveraging separate speaker and environment prompts, DAIEN-TTS allows independent control over the timbre and background environment of the synthesized speech. Built upon F5-TTS, the proposed DAIEN-TTS first incorporates a pretrained speech-environment separation (SES) module to disentangle the environmental speech into mel-spectrograms of clean speech and environment audio. Two random span masks of varying lengths are then applied to both mel-spectrograms, which, together with the text embedding, serve as conditions for infilling the masked environmental mel-spectrogram, enabling the simultaneous continuation of personalized speech and time-varying environmental audio. To further enhance controllability during inference, we adopt dual classifier-free guidance (DCFG) for the speech and environment components and introduce a signal-to-noise ratio (SNR) adaptation strategy to align the synthesized speech with the environment prompt. Experimental results demonstrate that DAIEN-TTS generates environmental personalized speech with high naturalness, strong speaker similarity, and high environmental fidelity.
@@ 2.3.2 SNR Adaptation
Accordingly, we can derive that $Scale=\sqrt{\|\boldsymbol{Y}^{S}_{spk}\|_{2}^{2}/\|\boldsymbol{Y}^{S}_{env}\|_{2}^{2}}$. During inference, $\boldsymbol{Y}^{E}_{env}$ is scaled by the calculated factor and then transformed into a mel-spectrogram, which serves as the final environment condition for environmental speech generation.
@@ 3.1 Dataset and Experimental Setup
In the SES module, we set the number of Transformer layers $K$ to 8, with 16 attention heads, an embedding dimension of 1024, and a feed-forward network (FFN) dimension of 2048. In the TTS module, each DiT block is augmented with a cross-attention layer, also configured with 16 attention heads. During training, we adopted a dynamic mixing strategy to construct clean–environmental speech pairs. For pretraining the SES module, all speech data was mixed with environment audio to ensure robust separation learning. For TTS training, we applied a probabilistic augmentation strategy, where each speech sample was mixed with environmental audio with a probability of 50%, and with silent segments otherwise, in order to encourage better alignment learning between text and speech. Both the SES module and the TTS module were trained for 600k steps with a batch size of 102,800 audio frames on 24 NVIDIA V100 32G GPUs, following the same training strategy as F5-TTS. For inference, we set the DCFG strength $\alpha_{speech}$ and $\alpha_{env}$ to 2.0. Audio samples can be accessed at https://yxlu-0102.github.io/DAIEN-TTS.
@@ 4.1 Evaluation with Silence Environment Prompts
The results of environment-aware TTS with silence environment prompts are shown in the upper part of Table 1. We first took the human recordings from the Seed-TTS test-en set as the ground truth (denoted as “Human”), and resynthesized them with the vocoder (denoted as “Vocoder”). Experimental results demonstrated that the vocoder resynthesized ground truth exhibited slight degradations in SIM-o and WER, which can serve as the upper bound for clean synthesized speech in terms of objective metrics. For the F5-TTS baseline trained on clean speech data, the synthesized speech exhibited high naturalness, intelligibility, and speaker similarity when provided clean speaker prompts. However, a noticeable drop in both objective and subjective metrics was observed when environmental speaker prompts were conditioned, indicating that F5-TTS was sensitive to background environment.